a potentiometer of wire length 1.0m has a resistance of 15ohms .it is connected to a 5volt battery in series with a resistance of 5ohms .determine the emf of the primary cell which gives a balance point at 60cm.
1
Expert's answer
2019-02-17T09:17:11-0500

According to Ohm's law the current is:


"I=\frac{E}{R+r}=\frac{5}{15+5}=0.25\text{ A}."


Voltage drop across the potential wire according to the same law:


"V=IR=0.25\cdot15=3.75\text{ V}."


Potential gradient on the wire:


"\sigma=V/L=3.75/1=3.75\text{ V/m}."


Electromotive force of the primary cell:


"E_\text{p}=\sigma\cdot x=3.75\cdot 0.6=2.25\text{ V}."